answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How do you subtract standard deviations?

Square the standard deviations, subtract/add them and calculate the square root of the subtraction/sum.

StDV=sqrt (StDvA^2+StDvB^2)

If you throw a coin 6 times how many possible outcomes are there?

Whenever you are trying to figure out the answer to an outcome problem, you just multiply how many sides it has by how many times you are tossing the coin.... 2 x 6 = 12 times.

===================================

Very reasonable. Warm, fuzzy, and intuitively satisfying. But, sadly, wrong.

Every toss of a coin has 2 possible outcomes.

If you write down the results of 6 tosses like: H T T H T H with an 'H' for each heads

and a 'T' for each tails, the number of different patterns you could write down for

six tosses is

2 x 2 x 2 x 2 x 2 x 2 = 64 .

If you don't care about the sequence, and you just want to know how many

different splits of 'heads' vs 'tails' you could get in 6 tosses, then there are seven

different possible outcomes:

-- no heads, 6 tails

-- 1 heads, 5 tails

-- 2 heads, 4 tails

-- 3 heads, 3 tails

-- 4 heads, 2 tails

-- 5 heads, 1 tails

-- 6 heads, no tails

What is the probability of rolling a sum of 10 when you roll two dice?

Each cube can land in 6 different ways.

Two cubes can land in (6 x 6) = 36 different ways.

There are 3 different ways of rolling a ten:

4 - 6

5 - 5

6 - 4

The probability of rolling a ten is (3 / 36) = 1/12 = 8 and 1/3 percent

What does probability have to do with heredity?

There is a chance of what u will recieve from ur parents genes.

Is 1.001 a probability?

No 1.001 is not a probability. Probability can not be >1

Mean and standard deviation of probability density function?

Suppose the probability density function is f(x), defined over a domain D

Then the mean is E(X) = x*f(x) integrated with respect to x over D.

Calculate E(X2) = x2*f(x) integrated with respect to x over D.

Then Variance(X) = E(X2) - [E(X)]2

and Standard Deviation = sqrt(Variance).

9 In how many different ways can you arrange 7 books on a shelf?

Consider one placement at a time. The first book on the shelf could be any of seven. The second could be any of the remaining six. This continues with the rest. The total combinations are:

7! = 7*6*5*4*3*2*1 = 5040

If two dice are thrown what is the probability that at least one of them shows a number greater than 3?

This is the same as 1 minus the probability that neither of them are greater than three. This is 1, minus the probability of getting greater than three, squared. Rolling higher than three has a 1/2 probability, so:

P(at least one greater than 3) = 1 - (1/2)2 = 1 - 1/4 = 3/4

If the number of customers follows a Poisson distribution with a rate of 5.6 per minute what is the distribution of the time it will take for the next two customers to arrive?

Average number of arriving customers in 4 minute interval is 2.8 and poison distributed what is the probability that exactly six customer will arrive in a 4 minute interval?

What is the sample size for a population of 60010?

well a sample size can be any size depending on the requirements. A sample size could be 10 people of that entire population or it could be 1000 people.

Trending Questions
What is the three forms of matter called? Jackie tossed a coin 25 times. Tails turned up 14 times. What was her experimental probability of gettting heads? How many 5 letter combinations can you make with the vowels? Why is there 1 dice that is red? What is the chance the 2 cards drawn from a stack of cards numbered 1 through 10 will have a sum of 9? What is the probability of tossing two dice and showing at least one 4? What is the probability of getting heads when you toss a coin? How many arrangements can be made out of the letters of the word mother taken four at a time so that each arrangement contains the letter'm'? What is the z score for 96 percent? What is rate of return on 30 percent probability and 12 percent returns? How do you solve Pp or Ppk if standard deviation is 0? One fifth of seths 10 marbles are blue how many of the marbles are blue? What is the probability of rolling a pair of six-sided dice 25 times and getting double sixes once Give your answer as a decimal rounded to the nearest hundredth? What is the probability of drawing a red card or a spade in two draws without replacing the card? What is the probability that z is greater than .5? How many outcomes are possible for a family of 5 children? How many colors would be needed to color a map of US states without any state touching the same colors? How many ways can the letters in the word mississippi be rearranged if the two Ps must stay together? What is the probability that a family of four children has at least 1 boy? What is the sample space of tossing a coin 4 times?